# Time Makes Space: Emergence of Place Fields in Networks Encoding Temporally Continuous Sensory Experiences

### v1

@@ Abstract
The vertebrate hippocampus is believed to use recurrent connectivity in area CA3 to support episodic memory recall from partial cues. This brain area also contains place cells, whose location-selective firing fields implement maps supporting spatial memory. Here we show that place cells emerge in networks trained to remember temporally continuous sensory episodes. We model CA3 as a recurrent autoencoder that recalls and reconstructs sensory experiences from noisy and partially occluded observations by agents traversing simulated arenas. The agents move in realistic trajectories modeled from rodents and environments are modeled as continuously varying, high-dimensional, sensory experience maps (spatially smoothed Gaussian random fields). Training our autoencoder to accurately pattern-complete and reconstruct sensory experiences with a constraint on total activity causes spatially localized firing fields, i.e., place cells, to emerge in the encoding layer. The emergent place fields reproduce key aspects of hippocampal phenomenology: a) remapping (maintenance of and reversion to distinct learned maps in different environments), implemented via repositioning of experience manifolds in the network’s hidden layer, b) orthogonality of spatial representations in different arenas, c) robust place field emergence in differently shaped rooms, with single units showing multiple place fields in large or complex spaces, and d) slow representational drift of place fields. We argue that these results arise because continuous traversal of space makes sensory experience temporally continuous. We make testable predictions: a) rapidly changing sensory context will disrupt place fields, b) place fields will form even if recurrent connections are blocked, but reversion to previously learned representations upon remapping will be abolished, c) the dimension of temporally smooth experience sets the dimensionality of place fields, including during virtual navigation of abstract spaces.
@@ 1 Introduction
The hippocampus is a brain region that plays a critical role in both spatial navigation knierimHippocampus2015 ; moserPlaceCellsGrid2015 ; moserSpatialRepresentationHippocampal2017 and episodic memory squireHippocampusMemorySpace1991 ; corkinWhatNewAmnesic2002 ; squireLegacyPatientNeuroscience2009 ; hasselmoHowWeRemember2011 ; hasselmoThetaRhythmEncoding2014 . Researchers have therefore sought to understand the neural substrates of these forms of memory in the hippocampus, as well as the extent to which these roles are interrelated eichenbaumCanWeReconcile2014 ; hasselmoModelsSpatialTemporal2017 ; eichenbaumIntegrationSpaceTime2017 ; whittingtonTolmanEichenbaumMachineUnifying2019 . A key role is played by place cells – so named because they fire in specific spatial locations within an environment okeefeHippocampusSpatialMap1971 . Place cells have spatially constrained firing fields, and “remap” firing patterns in response to significant contextual changes (especially in hippocampal subregion CA3) mullerEffectsChangesEnvironment1987 ; bostockExperienceDependentModifications1991 ; jezekThetapacedFlickeringPlacecell2011 ; solstadPlaceCellRate2014 ; leutgebRemappingDiscriminateContexts2014 ; markThetaPacedFlickering2017 , including motion to a new environment okeefeHippocampusCognitiveMap1978 and modified behavioral context or sensory cues almePlaceCellsHippocampus2014 ; kubieHippocampalRemappingPhysiological2020 . After remapping, they form almost orthogonal representations almePlaceCellsHippocampus2014 ; posaniFunctionalConnectivityModels2017 . These findings suggest that CA3 place cell network may play a key role in maintaining both positional and contextual information markThetaPacedFlickering2017 ; posaniIntegrationMultiplexingPositional2018 . Meanwhile, others have suggested that hippocampal CA3 is a pattern completion and separation device marrSimpleMemoryTheory1971 ; yassaPatternSeparationHippocampus2011 ; rollsMechanismsPatternCompletion2013 ; knierimTrackingFlowHippocampal2016 that supports the storage and retrieval of episodic memories from partial cues rollsComputationalTheoryEpisodic2010 ; hasselmoHowWeRemember2011 ; knierimTrackingFlowHippocampal2016 . This claim is supported by the extensive recurrent collaterals in CA3 liHippocampalCA3Network1994 ; witterIntrinsicExtrinsicWiring2007 ; hasselmoHowWeRemember2011 , which could allow it to act as an autoassociative network retrieving a complete memory fron partial cues mcnaughtonHippocampalSynapticEnhancement1987 ; hasselmoEncodingRetrievalEpisodic1998 ; hasselmoHowWeRemember2011 ; knierimTrackingFlowHippocampal2016 .
@@
Prior studies reconciling the spatial and episodic memory perspectives on CA3 propose that hippocampus associates sensory cues with spatial information mizumoriPreservedSpatialCoding1989 ; goldRoleCA3Subregion2005 ; leutgebPatternSeparationDentate2007 ; leutgebPatternSeparationPattern2007 . In this view, given partial sensory signals, CA3 recalls and reconstructs associated spatial information. Further, Benna and Fusi bennaPlaceCellsMay2021 proposed the hippocampus as a memory compression device, contrasting and memorizing differences between multiple visits to a room, and showed that autoencoders with sparsity constraints develop place-like representations in the network encoding layer. Likewise raju2022space suggests that spatial awareness results from processing sequences of sensory inputs, and in a learning framework based on Hidden Markov Models, observed the emergence of place-like patterns. Furthermore, levenstein2024sequential trained an RNN on visual observations conditioned on the agent’s direction to predict near future observations, and also found localized spatial representations. However, several questions remain. What signals contribute to spatial information? How are memories from multiple visits compared? Do such frameworks reproduce the phenomenology of place cells, including remapping and reversion across rooms during realistic navigation to create orthogonal representations?
@@ 2 Results
2.1 Episodic memory during spatial traversal
We first construct a framework for modeling episodic memory, then predict behavior during navigation. First, assume that the hippocampus continuously receives inputs from other cortical areas such as the entorhinal cortex, prefrontal cortex, and thalamus stewardCellsOriginEntorhinal1976 ; fellemanDistributedHierarchicalProcessing1991 ; hasselmoHowWeRemember2011 ; knierimHippocampus2015 , collectively constituting a high-dimensional experience vector (EV), a point in an abstract space of possible experiences $\mathbb{E}$. An agent experiences this world through a temporal sequence of “events”. Each event produces a short sequence of EVs - a ‘segment’ of episodic memory. An single event of bounded duration should trace a brief, continuous trajectory in the experience space $\mathbb{E}$. Formally, let $\mathbf{E}=\{\mathbf{e}_{0},\mathbf{e}_{1},\ldots,\mathbf{e}_{T}\}$ represent the segment of episodic memory generated by a single event, where each $\mathbf{e}_{t}\in\mathbb{R}^{D}$ (for $t=0,1,\ldots,T$) denotes an EV at time $t$, $D$ is the dimension of the input experience space, and $T$ is the duration of this epsiode in discrete time steps. We take each time step to represent 50 ms in real time. The sequence is continuous such that for any $t$ and $t+1$, where $0\leq t<T$, there exists a non-zero correlation coefficient, $\rho(\mathbf{e}_{t},\mathbf{e}_{t+1})$. During spatial navigation, we expect the environment to be temporally stable, with location-dependent experiences. We therefore model EVs in a fixed room as weakly spatial modulated (WSM) signals (Fig. 1), defining a manifold in the experience space $\mathbb{E}$. In this view, episodic memories of traversal of an environment sample the WSM manifold that defines it.
Figure 1: a. (i) Example trajectory of an artificial agent in a 2D room in actual space $\mathbb{A}$. (ii) Each room is defined by a unique set of weakly spatial modulated (WSM) signals representing location-dependent sensory cues. Within a room, a WSM rate map is defined by $F=z\ast g(\sigma),F\in\mathbb{R}^{W\times H}$, where $z$ is a 2D Gaussian random field. $W$ and $H$ are the dimensions of the room. A room is defined by its WSM set, i.e., $R=[F_{1},F_{2},\cdots,F_{D}]^{T},R\in\mathbb{R}^{D\times W\times H}$. b. Training schematic of our RAE. An artificial agent explores room(s) defined by a unique set of WSM signals, as depicted in panel a. Agents receive location-specific sensory experience vectors $\mathbf{e}_{x,y}=R[:,x,y]$, where $\mathbf{e}_{x,y}\in\mathbb{R}^{D}$. The agent’s trajectory within a trial is thereby converted into a sequence of experience vectors. At every training step, we randomly sample $N_{batch}$ segments of $T_{s}$ seconds from episodic memories within a $T_{w}$ second window. These segments form a stack of memories used to train the RAE. Every EV in this stack is randomly masked to occlude between $r_{min}$ to $r_{max}$of the signal with added Gaussian noise $\epsilon$. The RAE is trained to reconstruct complete, noiseless experience vectors. The sampling window is shifted forward by $\Delta t$ after each step until the end of the trial.
2.2 Emergence of place-like patterns in a Recurrent Autoencoder (RAE)
During each trial, agents sampled from the unique set of WSM signals that defined that environment. (Fig. 1a.i & ii), while moving in trajectories that imitated rodents (random walks with low-probability changes of direction and speed, see Suppl. Sec. 1). We discretized each $T_{t}$ trial into $dt$ timesteps during which agents received location-specific local EVs with elements randomly masked to imitate factors limiting observation like occlusion and partial attention. We concatenated timesteps sampled from a decaying distribution (Suppl. Sec. 1.2) of recent experiences (Fig. 1b) into ‘episodic bouts’ and trained the RAE on these bouts. We added pre- and post-activation noise to the hidden layer to simulate variability inherent to sensory processing. The use of episodic bouts serves to encourage input/output projections to update in consideration of events from a more extended history. We use a mean-squared error (MSE) objective function for pattern-completion, and a MSE constraint on the total hidden layer firing rates to enforce realistic firing rates.
@@ 2.2 Emergence of place-like patterns in a Recurrent Autoencoder (RAE)
After a trial ends, we “record" from the network by letting the agent run freely in the trial room while gathering firing patterns of units within the network (Suppl. Sec. 1). During this recording block, we pause RAE updates, so no additional learning occurs. We analyze the recorded firing patterns in the same way as experimentally recorded place cells, by projecting onto 2D room maps for visualization (Suppl. Sec. 1.3) and calculating the so-called Spatial Information Content (SIC, see Suppl. Sec 2.1) brandonReductionThetaRhythm2011 . Fig. 2 displays firing profiles of 40 random hidden units, with their spatial information content indicated above. A significant number of cells display strong place-like firing patterns. For example, consider a typical choice of training and constraint parameters (Suppl. Sec. 1.2), for which $\sim$10of the hidden units are active with average firing rates above a threshold of 0.1 Hz. We find in these cases that $\sim$80of the active hidden units have SIC greater than 5 which is 20 times the SIC of the WSM inputs (0.29 $\pm$ 0.01; mean $\pm$ SD). Thus, place-like cells emerge from dynamics of the task and network structure. The parameters we used in our experiments are in Suppl. Sec. 1.2, while we also verified that their precise values do not impact the emergence of place fields (Suppl. Sec. 4.2).
@@ 2.3 Place field emergence reflects temporally continuous changes in sensory experiences
First consider neurons without recurrent connections. Given $n$-dimensional inputs, each neuron’s input projection and ReLU non-linear firing rate collectively define an activation boundary (AB), an $n-1$ dimensional plane dividing the experience space $\mathbb{E}$. This boundary occurs at input loci where the neuron transitions from inactive to active. More generally, the distance of an input $\mathbf{e}$ to a neuron’s activation boundary determines the response firing rate, regardless of the activation function. For example, in Fig. 3b, two 1D EMs, $E1$ and $E2$, lie within a 2D experience space and are encoded by two neurons, $N1$ and $N2$, with the indicated ABs (solid blue for $N1$; solid orange for $N2$). As an agent moves along $E1$, it intersects $N1$’s activation boundary. Thus, $N1$’s firing rate will gradually increase then decrease. In contrast, since $E1$ is roughly perpendicular to $N2$’s activation boundary, motion along $E1$ will change neuron $N2$’s firing rate monotonically.
@@
Our autoencoder linearly decodes sensory experience from the hidden unit activations as $\mathbf{\hat{e}}=\sum_{j\in S}r_{j}{\bf W}^{out}_{j}$ where $j$ indexes hidden units, ${\bf W}^{out}_{j}$ is the output weight vector and $S$ are units that fire at the given location. If the firing rate of a unit $i$ increases monotonically as an agent moves along the EM, the angle between $\mathbf{\hat{e}}$ and the experience encoded by unit $i$, $\mathbf{W}^{out}_{i}$, will decrease monotonically as this unit dominates. This will cause the reconstructed experience at many locations to be dominated by a single vector which cannot generally be correct. On the other hand, if the activation boundary for a unit is exactly parallel to an EM, it will have the same firing rate everywhere and will not contribute to discriminating locations. Thus, the most useful units for reconstructing sensory experience will have activation boundaries segmenting the EM like $N1$ in Fig. 3b relative to $E1$ or $N2$ relative to $E2$.
Figure 3: a. A room is defined by a unique set of WSM signals describing expected sensory experience at every location. The set of WSM signals converts a room to a hyperplane in experience space. b. Illustration of neurons N1 and N2 intersecting two experience manifolds E1 and E2 in a 2-dimensional experience space. The yellow and blue lines denote activation boundaries, where neurons transition from inactive (negative sign) to firing (positive sign).
@@
This argument implies that training should rotate activation boundaries of some hidden units to segment the generally nonlinear experience manifold (EM). To accommodate the scenario in Fig. 3b in networks without recurrent connections, neurons would require substantial rewiring of input projections $W^{in}$. Recurrence helps reallocate the contributions to reconstructed sensory experience, so that the largest contributions are made by units whose activation boundaries segment the EM. This switching of encoding neurons obviates the need for input projection rewiring.
@@
We can also see from this reasoning why place-like units develop in the hidden layers. A given hidden unit contributes $\mathbf{\hat{e}_{i}}=r_{i}W^{out}_{i}$ to the reconstructed experience. As we explained above, only hidden units whose activation boundaries segment the EM are useful for reconstructing experience, so learning will either rotate activation boundaries to segment the EM, or adapt output weights $W^{out}_{i}$ to remove the unit from the reconstruction. Consider the location of peak firing of a unit whose planar activation boundary intersects the curved EM as in Fig. 3b. Because of the continuity of the EM, adjacent locations will produce similar experiences. The region of space where the unit fires will correspond to the part of the EM that is segmented by the activation boundary (indicated by ‘$+$’ in Fig. 3b). Assuming that the component of the experience vector that is orthogonal to the activation boundary varies approximately isotropically in space, the intersection locus of the activation boundary with the EM will be approximately circular when projected back to the environment. Thus, a population of such units, each tuned to part of the temporally continuous variation of experience, will exhibit place field like activation regions.

### v2

@@ Abstract
The vertebrate hippocampus is believed to use recurrent connectivity in area CA3 to support episodic memory recall from partial cues. This brain area also contains place cells, whose location-selective firing fields implement maps supporting spatial memory. Here we show that place cells emerge in networks trained to remember temporally continuous sensory episodes. We model CA3 as a recurrent autoencoder that recalls and reconstructs sensory experiences from noisy and partially occluded observations by agents traversing simulated arenas. The agents move in realistic trajectories modeled from rodents and environments are modeled as continuously varying, high-dimensional, sensory experience maps (spatially smoothed Gaussian random fields). Training our autoencoder to accurately pattern-complete and reconstruct sensory experiences with a constraint on total activity causes spatially localized firing fields, i.e., place cells, to emerge in the encoding layer. The emergent place fields reproduce key aspects of hippocampal phenomenology: a) remapping (maintenance of and reversion to distinct learned maps in different environments), implemented via repositioning of experience manifolds in the network’s hidden layer, b) orthogonality of spatial representations in different arenas, c) robust place field emergence in differently shaped rooms, with single units showing multiple place fields in large or complex spaces, and d) slow representational drift of place fields. We argue that these results arise because continuous traversal of space makes sensory experience temporally continuous. We make testable predictions: a) rapidly changing sensory context will disrupt place fields, b) place fields will form even if recurrent connections are blocked, but reversion to previously learned representations upon remapping will be abolished, c) the dimension of temporally smooth experience sets the dimensionality of place fields, including during virtual navigation of abstract spaces. Code for our experiments is available at [^1] https://zhaozewang.github.io/projects/time_makes_space.
@@ 1 Introduction
The hippocampus is a brain region that plays a critical role in both spatial navigation knierimHippocampus2015 ; moserPlaceCellsGrid2015 ; moserSpatialRepresentationHippocampal2017 and episodic memory squireHippocampusMemorySpace1991 ; corkinWhatNewAmnesic2002 ; squireLegacyPatientNeuroscience2009 ; hasselmoHowWeRemember2011 ; hasselmoThetaRhythmEncoding2014 . Researchers have therefore sought to understand the neural substrates of these forms of memory in the hippocampus, as well as the extent to which these roles are interrelated eichenbaumCanWeReconcile2014 ; hasselmoModelsSpatialTemporal2017 ; eichenbaumIntegrationSpaceTime2017 ; whittingtonTolmanEichenbaumMachineUnifying2019 . A key role is played by place cells – so named because they fire in specific spatial locations within an environment okeefeHippocampusSpatialMap1971 . Place cells have spatially constrained firing fields, and “remap” firing patterns in response to significant contextual changes (especially in hippocampal subregion CA3) mullerEffectsChangesEnvironment1987 ; bostockExperienceDependentModifications1991 ; jezekThetapacedFlickeringPlacecell2011 ; solstadPlaceCellRate2014 ; leutgebRemappingDiscriminateContexts2014 ; markThetaPacedFlickering2017 , including motion to a new environment okeefeHippocampusCognitiveMap1978 and modified behavioral context or sensory cues almePlaceCellsHippocampus2014 ; kubieHippocampalRemappingPhysiological2020 . After remapping, they form almost orthogonal representations almePlaceCellsHippocampus2014 ; posaniFunctionalConnectivityModels2017 . These findings suggest that CA3 place cell network may play a key role in maintaining both positional and contextual information markThetaPacedFlickering2017 ; posaniIntegrationMultiplexingPositional2018 , while also supporting a large capacity for memory storage battistaCapacityResolution2020 . Meanwhile, others have suggested that hippocampal CA3 is a pattern completion and separation device marrSimpleMemoryTheory1971 ; yassaPatternSeparationHippocampus2011 ; rollsMechanismsPatternCompletion2013 ; knierimTrackingFlowHippocampal2016 that supports the storage and retrieval of episodic memories from partial cues rollsComputationalTheoryEpisodic2010 ; hasselmoHowWeRemember2011 ; knierimTrackingFlowHippocampal2016 . This claim is supported by the extensive recurrent collaterals in CA3 liHippocampalCA3Network1994 ; witterIntrinsicExtrinsicWiring2007 ; hasselmoHowWeRemember2011 , which could allow it to act as an autoassociative network retrieving a complete memory fron partial cues mcnaughtonHippocampalSynapticEnhancement1987 ; hasselmoEncodingRetrievalEpisodic1998 ; hasselmoHowWeRemember2011 ; knierimTrackingFlowHippocampal2016 .
@@
Prior studies reconciling the spatial and episodic memory perspectives on CA3 propose that hippocampus associates sensory cues with spatial information mizumoriPreservedSpatialCoding1989 ; goldRoleCA3Subregion2005 ; leutgebPatternSeparationDentate2007 ; leutgebPatternSeparationPattern2007 . In this view, given partial sensory signals, CA3 recalls and reconstructs associated spatial information. Benna and Fusi bennaPlaceCellsMay2021 proposed the hippocampus as a memory compression device, contrasting and memorizing differences between multiple visits to a room, and showed that autoencoders with sparsity constraints develop place-like representations in the network encoding layer. Likewise raju2022space suggests that spatial awareness results from processing sequences of sensory inputs, and in a learning framework based on Hidden Markov Models, observed the emergence of place-like patterns. Furthermore, levenstein2024sequential trained an RNN on visual observations conditioned on the agent’s direction to predict near-future observations and also found localized spatial representations. However, several questions remain. What signals contribute to spatial information? How are memories from multiple visits compared? Do such frameworks reproduce the phenomenology of place cells, including remapping and reversion across rooms during realistic navigation to create orthogonal representations?
@@
2 Method
2.1 Episodic memory during spatial traversal
We first construct a framework for modeling episodic memory, then predict behavior during navigation. First, assume that the hippocampus continuously receives inputs from other cortical areas such as the entorhinal cortex, prefrontal cortex, and thalamus stewardCellsOriginEntorhinal1976 ; fellemanDistributedHierarchicalProcessing1991 ; hasselmoHowWeRemember2011 ; knierimHippocampus2015 , collectively constituting a high-dimensional experience vector (EV), a point in an abstract space of possible experiences $\mathbb{E}$. An agent experiences this world through a temporal sequence of “events”. Each event produces a short sequence of EVs - a ‘segment’ of episodic memory. A single event of bounded duration should trace a brief, continuous trajectory in the experience space $\mathbb{E}$. Formally, let $\mathbf{E}=\{\mathbf{e}_{0},\mathbf{e}_{1},\ldots,\mathbf{e}_{T}\}$ represent the segment of episodic memory generated by a single event, where each $\mathbf{e}_{t}\in\mathbb{R}^{D}$ (for $t=0,1,\ldots,T$) denotes an EV at time $t$, $D$ is the dimension of the input experience space, and $T$ is the duration of this episode in discrete time steps. We take each time step to represent 50 ms in real-time. The sequence is continuous such that for any $t$ and $t+1$, where $0\leq t<T$, there exists a non-zero correlation coefficient, $\rho(\mathbf{e}_{t},\mathbf{e}_{t+1})$. During spatial navigation, we expect the environment to be temporally stable, with location-dependent experiences. We therefore model EVs in a fixed room as weakly spatial modulated (WSM) signals (Fig. 1), defining a manifold in the experience space $\mathbb{E}$. In this view, episodic memories of traversal of an environment sample the WSM manifold that defines it.
2.2 Recurrent Autoencoder (RAE) as a model of pattern completion and separation
Figure 1: a. (i) Example trajectory of an artificial agent in a 2D room in actual space $\mathbb{A}$. (ii) Each room is defined by a unique set of weakly spatial modulated (WSM) signals representing location-dependent sensory cues. Within a room, a WSM rate map is defined by $F=z\ast g(\sigma),F\in\mathbb{R}^{W\times H}$, where $z$ is a 2D Gaussian random field. $W$ and $H$ are the dimensions of the room. A room is defined by its WSM set, i.e., $R=[F_{1},F_{2},\cdots,F_{D}]^{\top},R\in\mathbb{R}^{D\times W\times H}$. b. Training schematic of our RAE. An artificial agent explores room(s) defined by a unique set of WSM signals, as depicted in panel a. Agents receive location-specific sensory experience vectors $\mathbf{e}_{x,y}=R[:,x,y]$, where $\mathbf{e}_{x,y}\in\mathbb{R}^{D}$. The agent’s trajectory within a trial is thereby converted into a sequence of experience vectors. At every training step, we randomly sample $N_{batch}$ segments of $T_{s}$ seconds from episodic memories within a $T_{w}$ second window. These segments form a stack of memories used to train the RAE. Every EV in this stack is randomly masked to occlude between $r_{min}$ to $r_{max}$of the signal with added Gaussian noise $\epsilon$. The RAE is trained to reconstruct complete, noiseless experience vectors. The sampling window is shifted forward by $\Delta t$ after each step until the end of the trial.
During each trial, agents sampled from the unique set of WSM signals that defined that environment. (Fig. 1a.i & ii), while moving in trajectories that imitated rodents (random walks with low-probability changes of direction and speed, see Suppl. Sec. 1). We discretized each $T_{t}$ trial into $dt$ timesteps during which agents received location-specific local EVs with elements randomly masked to imitate factors limiting observation like occlusion and partial attention. We concatenated timesteps sampled from a decaying distribution (Suppl. Sec. 1.2) of recent experiences (Fig. 1b) into ‘episodic bouts’ and trained the RAE on these bouts. We added pre- and post-activation noise to the hidden layer to simulate variability inherent to sensory processing. The use of episodic bouts serves to encourage input/output projections to update in consideration of events from a more extended history. We use a mean-squared error (MSE) objective function for pattern completion and an MSE constraint on the total hidden layer firing rates to enforce realistic firing rates.
After a trial ends, we “record" from the network by letting the agent run freely in the trial room while gathering firing patterns of units within the network (Suppl. Sec. 1). During this recording block, we pause RAE updates so no additional learning occurs. We analyze the recorded firing patterns in the same way as experimentally recorded place cells, by projecting onto 2D room maps for visualization (Suppl. Sec. 1.3) and calculating the Spatial Information Content (SIC, see Suppl. Sec 2.1) brandonReductionThetaRhythm2011 .
@@ 3 Results
3.1 Emergence of place-like patterns
@@
Fig. 2 displays firing profiles of 40 random hidden units, with their spatial information content indicated above. A significant number of cells display strong place-like firing patterns. For example, consider a typical choice of training and constraint parameters (Suppl. Sec. 1.2), for which $\sim$10of the hidden units are active with average firing rates above a threshold of 0.1 Hz. We find in these cases that $\sim$80of the active hidden units have SIC greater than 5, which is 20 times the SIC of the WSM inputs (0.29 $\pm$ 0.01; mean $\pm$ SD). Thus, place-like cells emerge from dynamics of the task and network structure. The parameters we used in our experiments are in Suppl. Sec. 1.2, while we also verified that their precise values do not impact the emergence of place fields (Suppl. Sec. 4.2).
@@ 3.2 Place field emergence reflects temporally continuous changes in sensory experiences
First consider neurons without recurrent connections. Given $n$-dimensional inputs, each neuron’s input projection and ReLU non-linear firing rate collectively define an activation boundary (AB), an $n-1$ dimensional plane dividing the experience space $\mathbb{E}$. This boundary occurs at input loci where the neuron transitions from inactive to active. More generally, the distance of an input $\mathbf{e}$ to a neuron’s activation boundary determines the response firing rate, regardless of the activation function. For example, in Fig. 3b & c, an animal traversing a 1D track traces an EM within a 2D experience space. The EM is encoded by two neurons, $N1$ (Fig. 3c.i) and $N2$ (Fig. 3c.iii), with the indicated ABs. As an animal moves from $loc1$ to $loc2$, its perceived experience moves along EM and intersects $N1$’s activation boundary. This motion causes $N1$’s firing rate to gradually increase then decrease (Fig. 3c.ii). In contrast, $N2$’s activation boundary is roughly perpendicular to the same EM, motion from $loc1$ and $loc2$ will change neuron $N2$’s firing rate monotonically (Fig. 3c.iv).
@@
Our autoencoder linearly decodes sensory experience from the hidden unit activations as $\mathbf{\hat{e}}=\sum_{j\in S}r_{j}{\bf W}^{out}_{j}$ where $j$ indexes hidden units, ${\bf W}^{out}_{j}$ is the output weight vector and $S$ are units that fire at the given location. If the firing rate of a unit $i$ increases monotonically as an agent moves along the EM, the angle between $\mathbf{\hat{e}}$ and the experience encoded by unit $i$, $\mathbf{W}^{out}_{i}$, will decrease monotonically as this unit dominates. This will cause the reconstructed experience at many locations to be dominated by a single vector which cannot generally be correct. On the other hand, if the activation boundary for a unit is exactly parallel to an EM, it will have the same firing rate everywhere and will not contribute to discriminating locations. Thus, the most useful units for reconstructing sensory experience will have activation boundaries segmenting the EM like $N1$ in Fig. 3c.i.
Figure 3: a. A room is defined by a unique set of WSM signals describing expected sensory experience at every location. The set of WSM signals converts a room to a hyperplane in experience space. b. Illustration of animal moving from $loc1$ to $loc2$ in a 1D track. c. i & iii) Illustration of two neurons $N1$ and $N2$ intersecting the experience manifold EM corresponds to the 1D track. ii & iv) Corresponding firing rates as the animal moves from $loc1$ to $loc2$. d. EM1 and EM2 are two experience manifolds corresponding to two rooms. The optimal encoding neuron for EM1 may be temporarily inhibited through recurrent connections when entering a new room, then reactivated upon returning to the previous room.
@@
This argument implies that training should rotate activation boundaries of some hidden units to segment the generally nonlinear experience manifold (EM). To simultaneously encode similar but differently positioned EMs (Fig 3d), in networks without recurrent connections, neurons would require substantial rewiring of input projections $W^{in}$. Recurrence helps reallocate the contributions to reconstructed sensory experience, so that the largest contributions are made by units whose activation boundaries segment the EM. This switching of encoding neurons obviates the need for input projection rewiring.
@@
We can also see from this reasoning why place-like units develop in the hidden layers. A given hidden unit contributes $\mathbf{\hat{e}_{i}}=r_{i}W^{out}_{i}$ to the reconstructed experience. As we explained above, only hidden units whose activation boundaries segment the EM are useful for reconstructing experience, so learning will either rotate activation boundaries to segment the EM, or adapt output weights $W^{out}_{i}$ to remove the unit from the reconstruction. Consider the location of peak firing of a unit whose planar activation boundary intersects the curved EM as in Fig. 3c.i. Because of the continuity of the EM, adjacent locations will produce similar experiences. The region of space where the unit fires will correspond to the part of the EM that is segmented by the activation boundary (indicated by ‘$+$’ in Fig. 3c & d). Assuming that the component of the experience vector that is orthogonal to the activation boundary varies approximately isotropically in space, the intersection locus of the activation boundary with the EM will be approximately circular when projected back to the environment. Thus, a population of such units, each tuned to part of the temporally continuous variation of experience, will exhibit place-field-like activation regions.